Find three consecutive even integers such that twice the largest is equal to two more than three times the smallest.
Answer by namra(1)

let
integer 1=x smallest =x
integer 2=y =x+2
integer 3=z largest =x+4
2(x+4)=3(x)+2
2x+8=3x+2
8-2=3x-2x
x=6
integer 1=x=6
integer 2=y=x+2=8
integer 3=z=x+4=10